Find the horizontal and vertical asymptotes of the following. Type “none” if the function does not have an asymptote.
f(x)= 2x-3/x^2+2
Answer:
Horizontal:
Vertical:

Answer by stanbon(48535)

Find the horizontal and vertical asymptotes of the following. Type “none” if the function does not have an asymptote.
f(x)= 2x-3/x^2+2
-------------
Horizontal:
y = 0x^2/1x^2 = 0/1 = 0
---------------
Vertical:
x^2 + 2 = 0
none
==================
Cheers,
Stan H.